OpenStudy (anonymous):

C or D? Darrien is building a model of the Globe Theatre stage using a scale of 1:36. On the model, the width of the stage is 15 inches. What is the actual width, in feet, of the stage? a. 2.5 feet b. 45 feet c. 180 feet d. 540 feet

OpenStudy (amistre64):

what does the scale tell us: 1 inch to 36 feet? 1 foot to 36 inches? or some other relationship?

OpenStudy (amistre64):

maybe its unitless, and its just 1 to 36 therefore 15inches times (1 to 36) = 15 inches to 15(36) inches

OpenStudy (amistre64):

my bet is on 36inches = 3 feet, and we are at 15(3) feet

OpenStudy (anonymous):

That's what confused me as well, the lack of specified units. At first I did 15(36) and got D, but when I saw that the 36 was probably inches, I divided 540 by 3 (3 feet in 36 inches) and got 180.

OpenStudy (amistre64):

15(36 inches) = 15(3feet) = 45 feet

OpenStudy (anonymous):

So because you're give 15ft you need to convert 36 inches to feet and then solve?

OpenStudy (amistre64):

if 1 inch is equal to 3 feet .. 15(1inch) = 15(3feet)

OpenStudy (anonymous):

Ah, I see now.

OpenStudy (amistre64):

im assuming the scale is 1inch of model is equal to 36inches of actual object

OpenStudy (anonymous):

That makes sense. Thank you! :)

OpenStudy (amistre64):

36/3 = 12 so dividing by 3 is not quite right :) 36/12 = 3 is better